OpenStudy (anonymous):

The length of a rectangle is four times its width. If the area of the rectangle is 100 ft2 , find its perimeter.

OpenStudy (anonymous):

l=4w 100=4w*w 100=4(w)square w=5 perimeter=2(l+w) p=2(25) p=50

OpenStudy (rushwr):

Area = width* length if the unknown width is "x" then the length is "4x" area = 100 = 4x *x \[4x ^{2} =100\] \[x ^{2} = 25\] \[x =5\] So the width is 5 and the length is 4*5 = 20 Perimeter = length *2 +width *2 permiter = (20 *2 ) + (5*2) = 40 +10 = 50